# Probing forced responses and causality in data-driven climate emulators: conceptual limitations and the role of reduced-order models

### v5

@@ Abstract
A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture both stationary statistics and responses to external perturbations. Current neural climate emulators aim to resolve the atmosphere–ocean system in all its complexity but often struggle to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments. To explore the origin of these limitations, we first examine a simplified dynamical system that retains key features of climate variability. We interpret the results through linear response theory, providing a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms. We argue that the ability of multiscale systems’ emulators to reproduce perturbed statistics depends critically on (i) identifying an appropriate coarse-grained representation and (ii) careful parameterizations of unresolved processes. For low-frequency climate dynamics, these insights highlight reduced-order models, tailored to specific processes and scales, as valuable alternatives to general-purpose emulators. We next consider a real-world application, developing a neural model to investigate the joint variability of the surface temperature field and radiative fluxes. The model infers a multiplicative noise process directly from data, largely reproduces the system’s probability distribution, and enables causal studies through forced responses. We discuss its limitations and outline directions for future work. These results expose key challenges in data-driven modeling of multiscale physical systems and underscore the value of coarse-grained, stochastic approaches, with response theory as a principled framework to guide model design.
@@ I Introduction
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [Pathak2022, gencast, Lam2023, Bi2023, fourcastnet3] and climate [Kochkov2023, Watt-Meyer2023, Watt-Meyer2024, Lucie, OceanNet, Bire, Samudra, Camulator, lucie3D, SamudraACE]. Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system. For example, [Samudra] and [Camulator] emulate the full ocean and atmosphere components of state-of-the-art climate models. Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons. They can potentially be used for tuning and sensitivity analysis [M2Lines], and to improve the study of rare events [Lancelin], offering a promising direction for fast and efficient tools to study climate. However, several limitations remain: for example, long rollouts can suffer from error accumulation that can lead to instabilities [Chris, PedramStability]; rare events are not captured if absent from the training set, limiting their applicability for rare-event analysis in a changing climate [PedramPNAS]; key atmospheric modes, such as the Quasi-Biennial Oscillation and the Southern Annular Mode, are not well represented [Baxter]. A limitation particularly relevant for climate science, as opposed to weather forecasting, is that high skill in reproducing stationary statistics does not ensure accurate responses in perturbation experiments, thereby limiting the use of such models for causal studies of climate dynamics. This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related. The work of Van Loon et al. [Senne] illustrates this limitation in the context of the “pattern effect” problem [blochjohnsonetal2024, Dong2019AttributingPacific], i.e. the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance. Using the ACE2 emulator [Watt-Meyer2024] trained on the ERA5 reanalysis [era5], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response. The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way. This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [Samudra, Camulator, ZhangACE]. Even if such issues are addressed in specific cases, as shown by Wu et al (2025) [Wu] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Such limitations are also likely to become more pronounced in coupled settings. Additionally, emulators’ skill in capturing forced responses is often assessed using single trajectories. In climate applications, however, the relevant quantities are probability distributions, particularly the responses of the mean and variance. We will show that focusing on the skill in capturing variance responses can reveal important limitations otherwise hidden when examining only the mean response. The limitations examined here are especially relevant for emulators built from observational data and intended to model “real-world” dynamics.
@@
  • Unknown equations, partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unobserved [e20070509], or even unknown [BaldovinEntropy]. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [NormalForms, Majda2010]. The model mimics the abstract structure of dynamical cores of large-scale climate models with a linear operator and energy conserving quadratic nonlinearities. It provides a simplified yet relevant mathematical framework to unambiguously explore fundamental (rather than technical) limitations in capturing forced responses in climate emulators. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response can be computed by convolving the forcing with an impulse response operator [MajdaBook, Marconi]. This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to general external perturbations. Therefore, linear response theory is here proposed as a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms [Baldovin]. In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system. In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [PENLAND1996534, Gammaitoni, Dalmedico], though typically without addressing responses to external perturbations. Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
@@ II.1 Reduced-order stochastic climate models: theoretical and practical considerations
In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g. temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales. Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained representation obtained via (i–iii). The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [MajdaPerspective, STROUNINE2010145], by retaining the first $n<N$ of such bases. The correspondent principal components (PCs), serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as principal interaction patterns [CROMMELIN], community detection [FabCausal] and autoencoders [latentPierre, PierreLatent] may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [HeldFDT, Pedram2]. Identifying the “right” coarse-grained representation depends on the specific goals of the researcher and remains one of the most difficult challenges in data-driven modeling, as no general mathematical principle exists to guide such choices from data alone in high-dimensional dynamical systems [FutureFromPast, Baldovin, Baldovin].
@@ II.2 Response theory: a (very) brief overview
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems requires to study the effect of interventions on the system [IsmaelNew]. The inferred response operator quantifies time-dependent cause-effect interactions across degrees of freedoms $x_{j}$ and $x_{k}$ [Aurell, Lucarini2018, Baldovin, Valerio2024, FabCausal]. We refer to Appendix A for details on the relation between the response operator and causality.
@@ III Limitations of data-driven modeling by analogy: the case of partial observations
A fundamental limitation of purely data-driven modeling of real-world, multiscale climate systems is that many degrees of freedom remain unobserved. Even when the governing equations are known, as in meteorology, climate fields are only sampled at finite temporal and spatial resolutions, which requires data-driven closures for unresolved fast and small-scale processes. The challenge is even greater in settings where governing equations are unknown and provide no guidance for emulator design. In both cases, data-driven approaches can reproduce the variability of the retained variables, including their probability distributions, autocorrelation functions, and short-term forecasts. However, the response of probability distributions to perturbations, particularly the mean and variance in realistic climate applications, can uncover structural deficiencies stemming from unresolved dynamics that would otherwise remain hidden from stationary statistics alone. The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
@@ III.2.4 Main takeaways
While the results shown in this Section are encouraging, they remain far from real-world scenarios. Emulators of climate models would require full access to all climate fields at all temporal and spatial scales, which is often impractical due to the enormous data demands and the additional challenges posed by the need for very long time series for model training [FutureFromPast]. More fundamentally, emulators built from observational data and aimed at capturing “real-world” dynamics typically have access only to partial observations of a complex, interconnected system.
@@ III.3 The case of unknown equations and a partially observed state vector
  • Choice of variable to model. The reduced-order model depends critically on the choice of the modeled variable. Not all components can be selected to satisfy criteria (i) and (ii) discussed above. The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [MajdaBook, FutureFromPast, Lacorata, BaldovinPlos]. In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
@@
We note that for the specific case of the triad model in Eq. (6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [NormalForms, Majda2010]; see discussion in Appendix B. Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al. (2008) in the context of quasi-geostrophic modeling [STROUNINE2010145].. Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems. In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [Cecconi]: even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. The most severe limitation arises from the choice of variables: (i) the model’s skill depends critically on which variables are observed and included. Once these variables are fixed, (ii) the influence of unobserved degrees of freedom must be adequately parametrized to capture responses in probability distributions. Although these issues can be illustrated with low-dimensional examples, they become much more severe in realistic spatiotemporal climate systems, where time-scale separation is weak and variable selection involves nontrivial decisions about which fields to include, how to perform dimensionality reduction, and which spatial and temporal scales to retain — that is, how to construct an appropriate coarse-grained representation of the system. There is no rigorous mathematical framework to guide these choices from data alone in high-dimensional dynamical systems.
These considerations highlight intrinsic limitations of general-purpose emulators in capturing responses in mean and variance of the distribution to external perturbations. In contrast, reduced-order neural models are often tailored to specific problems, allowing one to leverage prior literature, insights, and results to guide variables and coarse-graining choices. Such prior knowledge can be essential for constructing data-driven models and for clearly delineating their domains of validity.
@@ IV.2.1 Data, preprocessing and the choice of proper variables
We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals. Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [FDTPatternEffect]. This averaging filters out higher-frequency processes that mediate surface–TOA coupling, leading to a “more Markovian” system . Given so, direct causal links between TOA fluxes at SST patterns can be detected as one time step responses; see Appendix A. We then high-pass filter the data with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration. This choice sacrifices the ability to capture multidecadal variability but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size. Implicitly, the choice assumes that the response of the net radiative flux to SST perturbations occurs within a 10-year window. Next, we project the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ onto an empirical orthogonal functions (EOFs) basis:

### v6

@@ Abstract
A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture both stationary statistics and responses to external perturbations. Current neural climate emulators aim to resolve the atmosphere–ocean system in all its complexity but often struggle to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments. To explore the origin of these limitations, we first examine a simplified dynamical system that retains key features of climate variability. We interpret the results through linear response theory, providing a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms. We argue that the ability of emulators of multiscale systems to reproduce perturbed statistics depends critically on (i) the choice of an appropriate coarse-grained representation and (ii) careful parameterizations of unresolved processes. These insights highlight reduced-order models, tailored to specific goals, processes, and scales, as valuable alternatives to general-purpose emulators. We next consider a real-world application, developing a neural model to investigate the joint variability of the surface temperature field and radiative fluxes. The model infers a multiplicative noise process directly from data, largely reproduces the system’s probability distribution, and enables causal studies through forced responses. We discuss its limitations and outline directions for future work. These results expose key challenges in data-driven modeling of multiscale physical systems and underscore the value of coarse-grained, stochastic approaches, with response theory as a principled framework to guide model design and enhance causal understanding.
@@ I Introduction
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [Pathak2022, gencast, Lam2023, Bi2023, fourcastnet3] and climate [Kochkov2023, Watt-Meyer2023, Watt-Meyer2024, Lucie, OceanNet, Bire, Samudra, Camulator, lucie3D, SamudraACE]. Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system. For example, [Samudra] and [Camulator] emulate the full ocean and atmosphere components of state-of-the-art climate models. Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons. They can potentially be used for tuning and sensitivity analysis [M2Lines], and to improve the study of rare events [Lancelin], offering a promising direction for fast and efficient tools in climate modeling. However, several limitations remain: for example, long rollouts can suffer from error accumulation that can lead to instabilities [Chris, PedramStability]; rare events are not captured if absent from the training set, limiting their applicability for rare-event analysis in a changing climate [PedramPNAS]; key atmospheric modes, such as the Quasi-Biennial Oscillation and the Southern Annular Mode, are not well represented [Baxter]. A limitation particularly relevant for climate science, as opposed to weather forecasting, is that high skill in reproducing stationary statistics does not ensure accurate responses in perturbation experiments, thereby limiting the use of such models for causal studies of climate dynamics. This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related. The work of Van Loon et al. [Senne] illustrates this limitation in the context of the “pattern effect” problem [blochjohnsonetal2024, Dong2019AttributingPacific], i.e. the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance. Using the ACE2 emulator [Watt-Meyer2024] trained on the ERA5 reanalysis [era5], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response. The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way. This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [Samudra, Camulator, ZhangACE]. Even if such issues are addressed in specific cases, as shown by Wu et al (2025) [Wu] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Such limitations are also likely to become more pronounced in coupled settings. Additionally, emulators’ skill in capturing forced responses is often assessed using single trajectories. In climate applications, however, the relevant quantities are probability distributions, particularly the responses of the mean and variance. We will show that focusing on the skill in capturing variance responses can reveal important limitations otherwise hidden when examining only the mean response. The limitations examined here are especially relevant for emulators built from observational data and intended to model “real-world” dynamics.
@@
  • Unknown equations, partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unobserved [e20070509], for example because of unresolved fast and small-scale processes, or even unknown [BaldovinEntropy]. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [NormalForms, Majda2010]. The model mimics the abstract structure of dynamical cores of large-scale climate models with a linear operator and energy conserving quadratic nonlinearities. It provides a simplified yet relevant mathematical framework to unambiguously explore fundamental (rather than technical) limitations in capturing forced responses in climate emulators. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response can be computed by convolving the forcing with an impulse response operator [MajdaBook, Marconi]. This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to general external perturbations. Therefore, linear response theory is here proposed as a rigorous framework to evaluate neural models beyond stationary statistics and probe causal mechanisms [Baldovin]. In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system. In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes. We then note that the lack of clear scale separation in realistic applications further complicates the problem, leading to non-unique coarse-graining choices tailored to specific goals. These issues are illustrated through a real-world example in a later section. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [PENLAND1996534, Gammaitoni, Dalmedico], though typically without addressing responses to external perturbations. Therefore, we will argue that a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
@@ II.1 Reduced-order stochastic climate models: theoretical and practical considerations
In spatiotemporal dynamical systems like climate, coarse-graining involves: (i) choosing a limited set of climate fields (e.g. temperature, velocities etc.), (ii) retaining a few relevant modes or large-scale averages, (iii) restricting the range of temporal scales. Throughout the paper, we will often refer to the so-called proper variables as the coarse-grained representation obtained via (i–iii). The relevant modes are often identified through Empirical Orthogonal Functions (EOFs) [MajdaPerspective, STROUNINE2010145], by retaining the first $n<N$ of such bases. The correspondent principal components (PCs), serve as the reduced time series for model training. Although EOFs are convenient, they may not be always optimal, and alternatives such as principal interaction patterns [CROMMELIN], community detection [FabCausal] and autoencoders [latentPierre, PierreLatent] may be more suitable depending on the application. The performance of the final model is sensitive to all coarse-graining choices (i–iii); see, e.g., [HeldFDT, Pedram2]. In climate dynamics, there is no clear space- and time- scale separation. In this case, identifying the “appropriate” coarse-grained representation depends on the specific goals of the researcher and remains one of the most difficult challenges in data-driven modeling, as no general mathematical principle exists to guide such choices from data alone in high-dimensional systems [FutureFromPast, Baldovin, Baldovin].
@@ II.2 Response theory: a (very) brief overview
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems requires to study the effect of interventions on the system [IsmaelNew]. The inferred response operator quantifies time-dependent cause-effect interactions across degrees of freedoms $x_{j}$ and $x_{k}$ [Aurell, Lucarini2018, Baldovin, LucariniPRL, FabCausal]. We refer to Appendix A for details on the relation between the response operator and causality.
@@ III Limitations of data-driven modeling by analogy: the case of partial observations
A fundamental limitation of purely data-driven modeling of real-world, multiscale climate systems is that many degrees of freedom remain unobserved. Even when the governing equations are known, as in meteorology, climate fields are only sampled at finite temporal and spatial resolutions, leading to unresolved fast and small-scale processes. The challenge is even greater in settings where governing equations are unknown and provide no guidance for emulator design. In both cases, data-driven approaches can reproduce the variability of the retained variables, including their probability distributions, autocorrelation functions, and short-term forecasts. However, the response of probability distributions to perturbations, particularly the mean and variance in realistic climate applications, can uncover structural deficiencies stemming from unresolved dynamics that would otherwise remain hidden from stationary statistics alone. The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
@@ III.2.4 Main takeaways
While the results shown in this Section are encouraging, they remain far from real-world scenarios. Emulators of climate models would require full access to all climate fields at all temporal and spatial scales, which is often impractical due to the enormous data demands. However, if that requirement were met, then the main challenge would reduce to the need for very long time series for model training [FutureFromPast]. More fundamentally, emulators built from observational data and aimed at capturing “real-world” dynamics typically have access only to partial observations of a complex, interconnected system.
@@ III.3 The case of unknown equations and a partially observed state vector
  • Choice of variable to model. The reduced-order model depends critically on the choice of the modeled variable. Not all components can be selected to satisfy criteria (i) and (ii) discussed above. The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [MajdaBook, FutureFromPast, Lacorata, BaldovinPlos]. In realistic spatiotemporal climate dynamics, such choice translates into selecting appropriate climate fields together with a narrow range of spatiotemporal scales. We will give a practical example in Section IV.
@@
Two brief remarks. (1) In partially observed settings, one might be tempted to appeal to Takens’ embedding theorem [Takens]. This is often not a practical option in real-world systems: the theorem does not extend to stochastic systems, and it becomes infeasible in high-dimensional settings [Cecconi, Baldovin, Lucente]. (2) For the specific case of the triad model in Eq. (6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [NormalForms, Majda2010]; see discussion in Appendix B. Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al. (2008) in the context of quasi-geostrophic modeling [STROUNINE2010145].. Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems. In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [Cecconi]: even before specifying a model structure, identifying the slow variables is nontrivial, non-unique and highly problem-dependent.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling multiscale real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. The most severe limitation arises from the choice of variables: (i) the model’s skill depends critically on which variables are observed and included. Once these variables are fixed, (ii) the influence of unobserved degrees of freedom must be adequately parametrized to capture responses in probability distributions. Although these issues can be illustrated with low-dimensional examples, they become much more severe in realistic spatiotemporal climate systems, where there is little or no scale separation and variable selection involves nontrivial, non-unique decisions about which fields to include, and which spatial and temporal scales to retain — that is, how to construct an appropriate coarse-grained representation of the system.
These considerations highlight intrinsic limitations of general-purpose emulators of coupled climate systems in capturing responses in mean and variance of the distribution to external perturbations. In contrast, reduced-order neural models are often tailored to specific problems, allowing one to leverage prior literature, insights, and results to guide variables and coarse-graining choices. Such prior knowledge can be essential for constructing data-driven models and for clearly delineating their domains of validity.
@@ IV.2.1 Data, preprocessing and the choice of proper variables
We coarse-grain the GFDL-CM4 control run in space and time, averaging daily data to monthly means and removing the seasonal cycle to focus on anomalies rather than periodic signals. Considering 1 month averages in TOA fluxes is not arbitrary: the immediate response of TOA fluxes to perturbations in SST is fast (e.g., $\sim 20$ days in tropical large-scale convective overturning motion in the GFDL-CM4 model) [FDTPatternEffect]. This averaging filters out higher-frequency processes that mediate surface–TOA coupling, leading to a “more Markovian” system . Given so, direct causal links between TOA fluxes and SST patterns can be detected as one time step responses; see Appendix A. We then high-pass filter the data with a cut-off of $10^{-1}$ years, removing (i) low-frequency (multidecadal) oscillations, which are poorly sampled even in 600-year runs, and (ii) slow drifts from incomplete ocean equilibration. This choice sacrifices the ability to capture multidecadal variability but ensures robust sampling of subdecadal statistics, yielding a larger effective sample size. Implicitly, the choice assumes that the response of the net radiative flux to SST perturbations occurs within a 10-year window. Next, we project the SST field $\mathbf{X}(t)\in\mathbb{R}^{N}$ onto an empirical orthogonal functions (EOFs) basis: